The area of a rectangle is 96 square ft. The two dimensions are in the ratio of two- thirds to one. Find the dimensions
Answer by issacodegard(60)

Hi, the very first thing to do is to characterize the problem mathematically.
So, I'll let the two dimensions be x and y.
Then we have,
xy=Area=96, and (2/3)x=y
So, by substitution we have,
x((2/3)x)=96, and so x^2=96(3/2)=144, so x=12
Then, the other dimension is y=(2/3)(12)=8
So the dimensions are 12 ft and 8 ft.